# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

### v2

@@ Introduction
Recent years have witnessed remarkable advancements in large language models (LLMs), resulting in a profound impact on various natural language processing (NLP) tasks 1, 2. These models have demonstrated unprecedented capabilities in generating and understanding human-like text, achieving outstanding performance across NLP tasks, such as summarization, question answering, and information retrieval. The exceptional performance of LLMs in core NLP tasks has catalyzed their exploration in the biomedical domain, where they show promise in supporting clinical decision-making, enhancing patient care quality, and improving clinical outcomes 3.
@@
However, the application of LLMs in the biomedical domain faces two critical challenges. First, these models often generate plausible-sounding but factually incorrect responses, a phenomenon commonly known as hallucination 4. This is particularly concerning in medical applications, where accuracy is paramount for patient safety. Second, the static nature of LLM training means that once the training process is complete, the model parameters remain fixed, resulting in an inability to access or incorporate up-to-date medical knowledge 5. This limitation is especially problematic in the rapidly evolving field of medicine, where new research findings and treatment guidelines emerge regularly.
@@
Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these critical challenges 6. By combining the generative capabilities of LLMs with dynamic access to external knowledge from trusted sources, RAG systems can: (1) provide transparent rationales for generated responses by explicitly referencing source documents; (2) maintain access to current biomedical knowledge through real-time information retrieval; (3) enhance the accuracy and reliability of medical information processing; and (4) reduce hallucination by grounding responses in the verified medical literature 1, 2, 5.
@@
The implementation of RAG in the biomedical domain presents unique challenges and opportunities. Biomedical information is characterized by complex terminology, specialized knowledge structures, and strict requirements for accuracy and reliability 7. Furthermore, the integration of RAG systems into clinical workflows requires careful consideration of factors such as information privacy, regulatory compliance, and clinical validation 8, 9.
Concurrent to our survey, several related surveys focused on biomedical-related LLMs. For example, Liu et al. 10 review medical LLMs and He et al. 2 discuss LLMs for healthcare. Similarly, Xiao 5 conducted a comprehensive overview of LLMs in medicine. Other discussed RAG systems in general domains, such as Fan et al. 1 and Gao et al. 11 focused on the architectures, training strategies, and application of RAG in general domains.
Our work differs from existing surveys in concentrating on the application of RAG in the biomedical domain, systematically reviewing its components, datasets, and challenges. Overall, this survey provides the following key contributions:
@@
    A systematic analysis of RAG technologies specifically adapted for biomedical applications, including specialized retrieval methods, ranking algorithms, and generation models.
@@
    A detailed examination of medical datasets, considering both knowledge source datasets and medical task datasets.
@@
    An exploration of current clinical applications and future directions, with particular attention to practical implementation challenges and emerging opportunities.
@@
    A structured framework for understanding the interplay between retrieval mechanisms, medical knowledge sources, and language models in biomedical RAG systems.
The remainder of this survey is organized as follows: Section 2 presents the background and the basic framework of biomedical RAG systems. Section 3 details the technical components, including retrieval methods, ranking algorithms, and generation models. Section 4 examines available datasets. Section 5 explores current clinical applications, and Section 6 discusses challenges and future directions.
Background
LLMs have disrupted NLP tasks with their ability to generate human-like text across diverse tasks 12. Despite their impressive capabilities, LLMs often struggle with knowledge limitations, hallucinations, and outdated information, particularly in specialized domains requiring precise information 13. RAG addresses these limitations by dynamically retrieving relevant external knowledge to supplement the LLM’s parametric knowledge before generating responses 11. This approach has shown particular promise in domain-specific applications, such as biomedicine, where accuracy, recency, and specialized knowledge are critical for clinical decision support, research analysis, and patient care 5.
In a typical workflow, a query first passes through the retriever, which searches relevant documents from knowledge sources. The re-ranking method then sorts these references based on their relevance. Finally, the top-k most relevant references, combined with the original query, are processed by the LLM to generate the response. As Fig. 1 shows, the basic framework of biomedical RAG consists of four key components 1: (1) Retriever: Responsible for indexing and retrieving relevant documents from various knowledge sources; Given a query $q$ and a collection of knowledge sources $K$, a retriever function $R(q,K)$ outputs a set of relevant documents $D=\{d_{1},d_{2},\ldots,d_{n}\}$, where each $d_{i}\in K$; (2) Knowledge Source: $K$ represents a heterogeneous collection of structured and unstructured data, including but not limited to medical literature, clinical guidelines, and databases; (3) Re-ranking Method:Re-sorting and prioritizing retrieved references, A re-ranking function $S(q,D)$ takes the query $q$ and the retrieved document set $D$ as input and outputs a re-ranked list of documents $D^{\prime}=[d_{1}^{\prime},d_{2}^{\prime},\ldots,d_{m}^{\prime}]$, where $D^{\prime}\subseteq D$ and the order reflects relevance to $q$; and (4) LLM: A LLM $G(q,D^{\prime})$ takes the original query $q$ and the re-ranked retrieved context $D^{\prime}$ as input to generate a final response $y$.
@@
RAG has been applied to a variety of biomedical applications. In clinical decision support systems, RAG powers evidence-based question-answering platforms like Clinfo.ai 14, enhances specialized decision-making through expert-guided LLMs 15, and facilitates rare disease diagnosis 16. This technology also strengthens diagnostic prediction by connecting electronic health records (EHRs) with medical knowledge repositories 17 and advances genomic medicine through improved variant annotation and interpretation 18. In the realm of medical imaging, RAG enhances report generation by seamlessly incorporating relevant prior studies and clinical guidelines 19. Across these applications, RAG leverages diverse data sources—including health records 20, 21, scientific literature 22, clinical guidelines 23, and genomic databases 18—to generate evidence-based responses that significantly improve both clinical care and research outcomes. The growing body of evidence consistently demonstrates that RAG holds remarkable promise for advancing information management and decision support in the biomedical and healthcare domains.
@@
In this section, we explore the technological underpinnings of biomedical RAG systems, which integrate specialized components for retrieval, re-ranking, and generation to effectively process domain-specific information. Our examination of retrieval mechanisms encompasses a range of approaches, including sparse 24, 25 and dense retrieval methods 26, 27, LLM-based retrieval 4, 28, knowledge graph enhancement techniques 29, 30, and emerging multi-modal retrieval capabilities designed to handle the diverse data formats prevalent in biomedical contexts 31, 19, 32. We further analyze re-ranking methodologies that refine initial retrieval results to prioritize the most relevant documents 33, 34, help address the challenge of relevance and correctness in biomedical information retrieval, where accuracy directly impacts clinical decision-making and research outcomes. Finally, we discuss the generation component of biomedical RAG systems, comparing closed-source 35, 36, 37 and open-source models 38, 39, 40, 41 alongside domain adaptation techniques 6, 42 that enable these general-purpose language models to produce specialized biomedical content with appropriate terminology, and contextual understanding.
@@
Retrievers are one of the core components of any RAG system, functioning as the mechanism that identifies and extracts relevant information from vast knowledge repositories in response to specific queries 11. In the biomedical domain, retrievers face distinctive challenges due to the heterogeneous nature of data structures—ranging from densely technical academic papers and specialized textbooks to complex biomedical knowledge graphs containing intricate entity relationships 43, 44. The domain’s specialized terminology, hierarchical concepts, and the life-critical nature of biomedical information make it more complex than information retrieval in general domains45. Effective biomedical retrievers must navigate these complexities while maintaining high precision to support accurate downstream analysis and generation. In this section, we describe various retrieval methodologies. We examine both standard retrieval approaches and how specific studies have modified these approaches for biomedical use, analyzing their respective strengths in handling domain-specific information structures and addressing unique retrieval challenges in the biomedical field.
@@
Sparse retrieval techniques represent a foundational approach in information retrieval, particularly valuable in biomedical RAG systems which rely on terminology precision and exact keyword matching. These methods, exemplified by algorithms like BM25 (Best Matching 25) 46 and TF-IDF (Term Frequency-Inverse Document Frequency) 47, operate on the principle of lexical matching, where documents are indexed based on term occurrences rather than semantic meaning. Despite the rise of dense retrieval methods, sparse retrievers maintain their relevance due to their interpretability, computational efficiency, and effectiveness, when dealing with highly specialized vocabulary that characterizes biomedical literature.
Table 1 shows some recent publications that utilize sparse retrieval methods in biomedical RAG systems. These papers demonstrate various applications of sparse retrieval, from pure BM25 48, 49 implementations to hybrid approaches 50, 51, 52, 53 that combine sparse and dense retrieval techniques. Notably, sparse retrieval is frequently treated as either a baseline for comparison or strategically incorporated as a component in ensemble solutions to increase retrieval diversity but also to improve generalization 54. Many of these works show that despite the attention given to newer neural approaches, traditional sparse retrieval methods remain competitive and often complementary to dense retrievers, particularly for specialized medical terminology and exact matching requirements 55, 56, 57. Their continued relevance in state-of-the-art biomedical RAG systems highlights the value of integrating lexical matching with semantic understanding when dealing with the precise terminology and complex nomenclature characteristic of biomedical literature.
In this line of work, BM25 stands out as the most widely adopted algorithm, featuring in more than half the studies (n=9) either independently or within hybrid systems. Established search tools including ElasticSearch 58, Whoosh 59, and Entrez API 60 also prove popular, offering accessible implementation options without requiring extensive customization. A notable trend is the increasing development of hybrid approaches that combine BM25 with dense retrieval techniques 61, 50, 51, 52, 53, leveraging both lexical precision and semantic understanding. The consistent appearance of sparse retrieval methods across all publication years, including the most recent 2025 studies, demonstrates their enduring value alongside newer neural approaches in cutting-edge biomedical information retrieval systems.
@@
Clinfo.ai 14 Entrez API Medical QA 2023
MEDRAG toolkit 61 Hybrid (BM25, SPECTER, Contriever, MedCPT) Medical QA 2024
SeRTS 56 Self-Rewarding Tree Search based BM25 Biomedical QA (BioASQ-QA) 2024
Dual RAG System 50 Hybrid (Dense + BM25 with language-specific tokenizers) Diabetes management 2024
MedExpQA 51 BM25 and MedCPT Medical Question Answering 2024
CliniqIR 52 BM25 and MedCPT Diagnostic decision support 2024
VAIV Bio-Discovery 53 Hybrid (Neural search + BM25) Biomedical knowledge discovery 2024
Genomics analysis GPT 18 Azure AI Search with keyword matching Genetic variant annotation 2025
Customized GPT 49 BM25 algorithm Medical QA 2025
Two-Layer RAG 48 Whoosh (Okapi BM25F ranking) Medical question answering 2025
LITURAt 62 Entrez API for PubMed retrieval Scientific data analysis 2025
@@
Dense retrievers represent a powerful approach to information retrieval that encodes queries and documents into continuous vector spaces to enable semantic matching based on learned representations rather than lexical overlap 63. These retrievers have become important in the biomedical domain where terminology is complex and specialized, often requiring a nuanced understanding of context and meaning beyond simple keyword matching 45. For instance, synonyms like ‘cancer’ and ‘neoplasm’ refer to the same concept, while hierarchical relationships exist between terms such as ’infectious disease’ and its specific type ’COVID-19’, further complicating accurate information retrieval. To analyze biomedical RAG systems, we categorize dense retrievers into five distinct types: Type 1: Pre-trained language models (PLMs); Type 2: Commercial embedding APIs; Type 3: Domain-adapted PLMs; Type 4: LLM-based embeddings; and Type 5: Domain-specific retrievers. In the following paragraphs, we provide a detailed analysis of each retriever type’s characteristics, followed by a comprehensive review of recent works that implement these approaches in biomedical information retrieval systems.
Type 1: PLMs PLMs are trained on general corpora and are widely used in various information retrieval systems to create dense representations of queries and documents, so-called embeddings. There are numerous open-source general embedding models available, which effectively capture semantic relationships and enable meaningful text matching, making them common foundational components in biomedical RAG systems. As shown in Table LABEL:tab:Classification, transformer-based PLMs like BERT, SBERT 64, and Sentence Transformer 65 have been widely adopted in recent biomedical retrieval systems. The continued popularity of these models is evident from numerous recent implementations. Additionally, newer embedding models such as BGE-embedding 66, Dense Passage Retriever 67 and GIST Large Embedding 68 have gained significant adoption in recent literature, further diversifying the range of options available for biomedical retrieval tasks. Several other widely-used PLMs, including FastText 69, RoBERTa 70, and SimCSE 71, though not specifically represented in Table LABEL:tab:Classification, remain viable alternatives for developing effective biomedical RAG applications.
@@
AskFDALabel 42 Sentence Transformer 65 FDA drug labeling extraction 2024
CaLM 26 BGE-embedding 66 with Chroma DB Supporting caregivers FM 2024
RAG-HPO 72 FastEmbed 73 Rare genetic disorders automated deep phenotyping 2024
RALL 27 Dense Passage Retriever 67 Lay language generation 2024
RT 74 BERT embeddings 75 Few-shot medical NER 2024
RAG-GPT 76 BGE-embedding 66 Breast cancer nursing care QA 2024
GNQA 77 HNSW graphs 78 Medical QA 2024
Klang et al. 79 GIST Large Embedding 68 with FAISS ICD-10-CM coding 2024
RAGPR 80 SBERT embeddings 64 Personalized physician recommendations 2025
JGCLLM 81 GLuCoSE-base-ja vectors 82 Genetic counseling support 2025

### v3

@@ Introduction
Recent years have witnessed remarkable advancements in large language models (LLMs), resulting in a profound impact on various natural language processing (NLP) tasks Fan2024RAG, he2023survey. These models have demonstrated unprecedented capabilities in generating and understanding human-like text, achieving outstanding performance across NLP tasks, such as summarization, question answering, and information retrieval. The exceptional performance of LLMs in core NLP tasks has catalyzed their exploration in the biomedicine domain, where they show promise in supporting clinical decision-making, enhancing patient care quality, and improving clinical outcomes Wang2024JMLR.
@@
However, the application of LLMs in the biomedical domain faces two critical challenges. First, these models often generate plausible-sounding but factually incorrect responses, a phenomenon commonly known as hallucination li2025biomedrag. This is particularly concerning in medical applications, where accuracy is paramount for patient safety. Second, the static nature of LLM training means that once the training process is complete, the model parameters remain fixed, resulting in an inability to access or incorporate up-to-date medical knowledge XIAO2025102888. This limitation is especially problematic in the rapidly evolving field of medicine, where new research findings and treatment guidelines emerge regularly.
Biomedicine RAG Dataset and Evaluation § IV Technology § III Preliminary § II Introduction § I Applications § V Discussion § VI Conclusion § VII Reranker Retriever Generation Domain Adaptation Medical Task Datasets Knowledge Source Datasets Evaluation Metrics Precision Medicine Applications Clinical Report Generation Clinical Decision Support Systems Healthcare Education Clinical Research Summary of dataset and evaluation approaches Summary of technological approaches Summary of RAG applications in biomedicine Challenges and Future Direction Knowledge Graph Enhanced Retrieval Dense Retrieval Sparse Retrieval Multi-modal Retrieval Commercial LLMs Open-Source LLMs Biomedical Specialized LLMs Retriever Stage Adaptation Reranker Stage Adaptation LLM Stage Adaptation Diagnostic and Treatment Decision Support Medical Question Answering Rare Disease Identification and Management Genomic Medicine Personalized Health Management
Figure 1: Structure of this paper
@@
Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these critical challenges frisoni2024generate. By combining the generative capabilities of LLMs with dynamic access to external knowledge from trusted sources, RAG systems can: (1) provide transparent rationales for generated responses by explicitly referencing source documents; (2) maintain access to current biomedical knowledge through real-time information retrieval; (3) reduce hallucinations by grounding responses in verified medical literature, thereby enhancing the accuracy and reliability of medical information processingFan2024RAG, he2023survey, XIAO2025102888.
@@
The implementation of RAG in the biomedical domain presents unique challenges and opportunities. Biomedical information is characterized by complex terminology, specialized knowledge structures, and strict requirements for accuracy and reliability xiong2024benchmarking. Furthermore, the integration of RAG systems into clinical workflows requires careful consideration of factors such as information privacy, regulatory compliance, and clinical validation 10.1007/978-981-97-9437-9_2, santra2024regulatory. Therefore, a systematic survey of existing studies applying RAG in the biomedicine and healthcare domain is necessary to understand how these challenges are being addressed and to identify current research trends and gaps.
We considered papers that are directly related to the application of RAG techniques in the biomedicine or healthcare domain. Eligible studies include those that incorporate RAG components—such as retrievers, rerankers(optional), and generation models—or apply RAG to improve model performance on biomedical tasks. Specifically, we included works applying RAG to biomedical named entity recognition (NER), biomedical question answering (QA), multiple-choice reasoning, diagnostic classification, and medical dialogue systems, including, but not limited to, these areas. Studies were required to demonstrate how RAG integration enhances the performance, accuracy, reliability, or factual grounding of biomedical models, systems, or applications. Studies that use RAG only as a baseline or comparative model without substantively integrating RAG into their methodological framework were excluded. To collect relevant literature, we systematically searched PubMed and Google Scholar for studies published between 2022 and 2025, ensuring coverage of the most recent developments in biomedical RAG research. In total, 143 papers met the inclusion criteria and were included in this survey.
Concurrent to our work, several surveys have examined either biomedical LLMs or general-domain RAG systems. Surveys such as Liu et al. liu2024survey and He et al. he2023survey focus primarily on medical LLMs, emphasizing model evolution, training data, clinical applications, and safety considerations, but they do not analyze the RAG pipeline or retrieval–generation integration. Xiao et al. XIAO2025102888 further expand to multimodal LLMs in medicine, yet similarly concentrate on model architectures and clinical use cases rather than grounding mechanisms or retrieval-oriented methodologies.
In contrast, recent surveys on RAG—such as Fan et al. Fan2024RAG and Gao et al. Gao2023RAGSurvey—provide systematic overviews of RAG architectures, training paradigms, and evaluation practices in general domains. However, they do not address the unique challenges of biomedical knowledge sources, medical-language characteristics, domain adaptation strategies, or multimodal medical data.
Our survey bridges this gap by offering the first retrieval-centric and pipeline-level review of RAG techniques specifically tailored for biomedical applications. We synthesize advances from sparse/dense/LLM-based retrieval, biomedical rerankers, and domain-adapted generation models, and further integrate developments in knowledge graphs and multimodal retrieval. By unifying these threads, our work positions biomedical RAG as a distinct research area and provides a structured roadmap that helps developers and researchers understand current progress and effectively contribute to future advances. Table 1 provides a brief comparison of recent related survey.
Table 1: Comparative summary of related survey papers
Reference Year RAG Biomedical Pipeline Coverage Key Contributions
He et al. he2023survey 2023 Generation only Provides an overview of medical LLM development, applications, and ethical considerations, but does not address retrieval or grounding mechanisms.
Liu et al. liu2024survey 2024 Generation only Covers Med-LLM technologies, applications, and trustworthiness, yet lacks discussion of retrieval augmentation or biomedical evidence integration.
Xiao et al. XIAO2025102888 2025 Generation only Reviews LLMs and MLLMs in medicine, emphasizing model architecture and clinical scenarios, but does not examine retrieval, reranking, or grounding.
Fan et al. Fan2024RAG 2024 Retrieval + Generation Provides a structured overview of RAG architectures and training strategies in general domains, without biomedical datasets or domain-specific methods.
Gao et al. Gao2023RAGSurvey 2023 Retrieval + Generation Introduces Naive, Advanced, and Modular RAG frameworks, but does not cover biomedical tasks, knowledge sources, or clinical evaluation.
Ours 2025 Full pipeline: retrieval, reranking, generation, domain adaptation Offers the first domain-specific, pipeline-level survey of biomedical RAG, integrating sparse/dense/LLM-based retrievers, biomedical rerankers, domain-adapted generators, knowledge-graph RAG, multimodal retrieval, and medically grounded evaluation practices.
Overall, this survey provides the following key contributions:
@@
    A comprehensive review of retrieval-augmented technologies tailored to the biomedical domain, covering sparse, dense, LLM-based, and domain-specialized retrievers, as well as biomedical rerankers and generation models.
@@
    A detailed examination of medical datasets and evaluation, considering both knowledge source datasets and medical task benchmarks, along with their evaluation methods.
@@
    A practical, pipeline-oriented framework that examines existing biomedical RAG applications and offers developers and practitioners a clear roadmap for system implementation, while also highlighting future challenges and research directions.
Fig. 1 illustrates the overall structure of this survey. The remainder of the paper is organized as follows. Section II introduces the background and fundamental framework of biomedical RAG systems. Section III details the core technical components, including retrieval methods, ranking algorithms, and generation models. Section IV reviews the available datasets and evaluation methodologies. Section V discusses biomedical applications that leverage RAG. Section VI summarizes the key observations and discusses current challenges and future directions. Finally, Section VII concludes the paper.
Preliminary
To provide a clear conceptual and structural understanding of biomedical RAG systems, we first summarize their general workflow and key components. As Fig. 2 shows, in a typical biomedical RAG workflow, the system can be formally characterized as a composition of sequential functions operating over a query and an external knowledge source. Given a user query $q$ and a heterogeneous knowledge collection $K$, the end-to-end RAG process generates a response $y$ through the functional pipeline:
$y=G\big(q,\,S\big(q,\,R(q,K)\big)\big),$ (1)
where each component $R$, $S$ and $G$ represents a distinct stage of the retrieval–generation paradigm.
First, the Retriever $R(q,K)$ identifies a set of candidate documents relevant to the query and returns a document collection $D=\{d_{1},d_{2},\ldots,d_{n}\}\subseteq K$. Next, the Re-ranking Method $S(q,D)$ is an optional component of the RAG pipeline that refines the retrieved set by sorting and prioritizing documents according to task-specific relevance criteria, producing a re-ranked subset $D^{\prime}=[d^{\prime}_{1},d^{\prime}_{2},\ldots,d^{\prime}_{m}]$ with $D^{\prime}\subseteq D$. The use of this component, as well as the choice of scoring strategy, varies across biomedical applications. Finally, the Generation Function $G(q,D^{\prime})$ synthesizes the retrieved evidence with the original query to produce the final output:
$y=G(q,D^{\prime})=\text{LLM}\big(\text{prompt}(q,D^{\prime})\big),$ (2)
where the prompting function integrates the query with the selected context, $\text{prompt}(q,D^{\prime})=q\oplus D^{\prime}$. This formulation describes RAG as a modular and extensible computational paradigm, whose behavior can be adapted to a wide range of biomedical tasks through different designs of the generation function $G$.
To illustrate this flexibility, we formalize 4 representative instantiations of $G$ in biomedical applications:
  1. 1.
    Biomedical Text Classification. For classification-oriented tasks, the generation function $G$ selects the most probable label from the label set $\mathcal{C}$ under the model’s conditional distribution:
    $G_{\text{cls}}(q,D^{\prime})=\arg\max_{c\in\mathcal{C}}\,p_{\theta}(c\mid\text{prompt}(q,D^{\prime})).$ (3)
    This formulation encompasses a variety of biomedical classification scenarios, including diagnostic classification and multiple-choice question answering.
  2. 2.
    Biomedical Question Answering. For free-text biomedical QA, $G$ generates an answer sequence $y=[y_{1},\ldots,y_{T}]$ by sampling (or decoding) tokens from the conditional language model distribution:
    $G_{\text{qa}}(q,D^{\prime})=\arg\max_{y}\,p_{\theta}(y\mid\text{prompt}(q,D^{\prime})),$ (4)
    where $y$ represents the generated natural-language answer.
  3. 3.
    Clinical Dialogue or Consultation Systems. For multi-turn dialogue, $G$ conditions on the dialogue history $h$ in addition to the retrieved evidence:
    $G_{\text{dialog}}(q,D^{\prime},h)=\arg\max_{y}\,p_{\theta}(y\mid h\oplus\text{prompt}(q,D^{\prime})),$ (5)
    where $y$ is the next system response.
@@
    Biomedical Named Entity Recognition (NER). For sequence labeling tasks, $G$ predicts a sequence of token-level labels:
    $G_{\text{ner}}(q,D^{\prime})=\big[\arg\max_{y_{t}\in\mathcal{L}}\,p_{\theta}(y_{t}\mid\text{prompt}(q,D^{\prime}),t)\big]_{t=1}^{T},$ (6)
    where $\mathcal{L}$ is the label set and $T$ is the input sequence length.
These formulations highlight how biomedical RAG systems form a unified retrieval–generation pipeline, while tailoring the generation function to meet the requirements of diverse downstream tasks.
Beyond the task-specific instantiations of the generation function $G$, the RAG paradigm further benefits from the inherent extensibility of its external knowledge source. The knowledge repository $K$ is designed to be continuously updatable, enabling biomedical RAG systems to maintain timely access to newly emerging medical evidence. As new documents or knowledge collections—such as updated clinical guidelines, recently published biomedical literature, or domain-specific databases—become available, they can be seamlessly incorporated into the existing repository. Formally, the expanded knowledge source is expressed as $K_{\text{updated}}=K\cup K_{\text{new}}$, where $K_{\text{new}}$ denotes the newly added knowledge. This extensibility ensures that retrieval and generation remain grounded in the most current biomedical information, thereby enhancing the robustness and adaptability of RAG across diverse clinical and research applications.
@@
RAG has been widely applied across diverse biomedical and healthcare scenarios, including clinical decision support, healthcare education, clinical report generation, and other related applications. Across these applications, RAG use diverse data sources-including health records johnson2016mimic, herrett2015cprd, scientific literature specter2020cohan, clinical guidelines Bashir2024, and genomic databases Lu2025BoostingGPT—to generate evidence-based responses that significantly improve both clinical care and research outcomes. A more detailed discussion of specific applications is provided in Section V.
@@
In this section, we explore the technological underpinnings of biomedical RAG systems, which integrate specialized components for retrieval, re-ranking, and generation to effectively process domain-specific information. Our examination of retrieval mechanisms encompasses a range of approaches, including sparse Jin2022, zhang2024online and dense retrieval methods parmanto2024calm, Guo2024, LLM-based retrieval li2025biomedrag, ong2024surgeryllm, knowledge graph enhancement techniques Kakalou2024, Ascle, and emerging multi-modal retrieval capabilities designed to handle the diverse data formats prevalent in biomedical contexts Sun2024FactAware, Raminedi2024MultimodalTransformer, Upadhyaya2024. We further analyze re-ranking methodologies that refine initial retrieval results to prioritize the most relevant documents jin2023medcpt, jia2024podgpt, help address the challenge of relevance and correctness in biomedical information retrieval, where accuracy directly impacts clinical decision-making and research outcomes. Finally, we discuss the generation component of biomedical RAG systems, comparing closed-source OpenAI2025TextGeneration, Anthropic2025GettingStarted, GoogleGemini2025 and open-source models MetaAI2024LLaMA3, MosaicML2023Introducing, mistral_small_24b_2024, deepseekr1_2024 alongside domain adaptation techniques frisoni2024generate, llm-regulatory-framework that enable these general-purpose language models to produce specialized biomedical content with appropriate terminology, and contextual understanding.
@@
Retrievers are one of the core components of any RAG system, functioning as the mechanism that identifies and extracts relevant information from vast knowledge repositories in response to specific queries Gao2023RAGSurvey. In the biomedical domain, retrievers face distinctive challenges due to the heterogeneous nature of data structures—ranging from densely technical academic papers and specialized textbooks to complex biomedical knowledge graphs containing intricate entity relationships 10.1145/3648471, 10.1145/3462476. The domain’s specialized terminology, hierarchical concepts, and the life-critical nature of biomedical information make it more complex than information retrieval in general domainssivarajkumar2024clinical. Effective biomedical retrievers must navigate these complexities while maintaining high precision to support accurate downstream analysis and generation. In this subsection, we present four main retrieval paradigms used in biomedical RAG systems. Sparse and dense retrieval represent the most commonly adopted approaches, whereas knowledge graph–enhanced and multimodal retrieval are designed to handle more complex data structures, such as structured knowledge graphs and multimodal biomedical information.
@@
Sparse retrieval techniques represent a foundational approach in information retrieval, particularly valuable in biomedical RAG systems which rely on terminology precision and exact keyword matching. These methods, exemplified by algorithms like BM25 (Best Matching 25) bm25 and TF-IDF (Term Frequency-Inverse Document Frequency) Jones1972, operate on the principle of lexical matching, where documents are indexed based on term occurrences rather than semantic meaning. Despite the rise of dense retrieval methods, sparse retrievers maintain their relevance due to their interpretability, computational efficiency, and effectiveness, when dealing with highly specialized vocabulary that characterizes biomedical literature.
Table 2 shows some recent publications that utilize sparse retrieval methods in biomedical RAG systems. These papers demonstrate various applications of sparse retrieval, from pure BM25 Das2025, Gong2024 implementations to hybrid approaches Lee2024, ALONSO2024102938, Abdullahi2024, Kim2024 that combine sparse and dense retrieval techniques. Notably, sparse retrieval is frequently treated as either a baseline for comparison or strategically incorporated as a component in ensemble solutions to increase retrieval diversity but also to improve generalization kamalloo:2024. Many of these works show that despite the attention given to newer neural approaches, traditional sparse retrieval methods remain competitive and often complementary to dense retrievers, particularly for specialized medical terminology and exact matching requirements xu2024bmretriever, Hu2024, izacard2021contriever. Their continued relevance in state-of-the-art biomedical RAG systems highlights the value of integrating lexical matching with semantic understanding when dealing with the precise terminology and complex nomenclature characteristic of biomedical literature.
In this line of work, BM25 stands out as the most widely adopted algorithm, featuring in more than half the studies (n=9) either independently or within hybrid systems. Established search tools including ElasticSearch ElasticsearchDocs, Whoosh WhooshDocs, and Entrez API EntrezAPI also prove popular, offering accessible implementation options without requiring extensive customization. A notable trend is the increasing development of hybrid approaches that combine BM25 with dense retrieval techniques xiong-etal-2024-benchmarking, Lee2024, ALONSO2024102938, Abdullahi2024, Kim2024, leveraging both lexical precision and semantic understanding. The consistent appearance of sparse retrieval methods across all publication years, including the most recent 2025 studies, demonstrates their enduring value alongside newer neural approaches in cutting-edge biomedical information retrieval systems.
@@
Clinfo.ai Lozano2024ClinfoAI Entrez API Medical QA 2023
MEDRAG toolkit xiong-etal-2024-benchmarking Hybrid (BM25, SPECTER, Contriever, MedCPT) Medical QA 2024
SeRTS Hu2024 Self-Rewarding Tree Search based BM25 Biomedical QA (BioASQ-QA) 2024
Dual RAG System Lee2024 Hybrid (Dense + BM25 with language-specific tokenizers) Diabetes management 2024
MedExpQA ALONSO2024102938 BM25 and MedCPT Medical Question Answering 2024
CliniqIR Abdullahi2024 BM25 and MedCPT Diagnostic decision support 2024
VAIV Bio-Discovery Kim2024 Hybrid (Neural search + BM25) Biomedical knowledge discovery 2024
Genomics analysis GPT Lu2025BoostingGPT Azure AI Search with keyword matching Genetic variant annotation 2025
Customized GPT Gong2024 BM25 algorithm Medical QA 2025
Two-Layer RAG Das2025 Whoosh (Okapi BM25F ranking) Medical question answering 2025
LITURAt Peasley2025 Entrez API for PubMed retrieval Scientific data analysis 2025
@@
Dense retrievers represent a powerful approach to information retrieval that encodes queries and documents into continuous vector spaces to enable semantic matching based on learned representations rather than lexical overlap 10.1145/3637870. These retrievers have become important in the biomedical domain where terminology is complex and specialized, often requiring a nuanced understanding of context and meaning beyond simple keyword matching sivarajkumar2024clinical. For instance, synonyms like ‘cancer’ and ‘neoplasm’ refer to the same concept, while hierarchical relationships exist between terms such as ’infectious disease’ and its specific type ’COVID-19’, further complicating accurate information retrieval. To analyze biomedical RAG systems, we categorize dense retrievers into five distinct types: Type 1: Pre-trained language models (PLMs); Type 2: Commercial embedding APIs; Type 3: Domain-adapted PLMs; Type 4: LLM-based embeddings; and Type 5: Domain-specific retrievers. In the following paragraphs, we provide a detailed analysis of each retriever type’s characteristics, followed by a comprehensive review of recent works that implement these approaches in biomedical RAG systems.
Type 1: PLMs. PLMs are trained on general corpora and are widely used in various information retrieval systems to create dense representations of queries and documents, so-called embeddings. There are numerous open-source general embedding models available, which effectively capture semantic relationships and enable meaningful text matching, making them common foundational components in biomedical RAG systems. As shown in Table LABEL:tab:Classification, transformer-based PLMs like BERT, SBERT santander2022semantic, and Sentence Transformer thakur2020augmented have been widely adopted in recent biomedical retrieval systems. The continued popularity of these models is evident from numerous recent implementations. Additionally, newer embedding models such as BGE-embedding bge_embedding, Dense Passage Retriever Karpukhin2020 and GIST Large Embedding solatorio2024gistembed have gained significant adoption in recent literature, further diversifying the range of options available for biomedical retrieval tasks. Several other widely-used PLMs, including FastText bojanowski2017enriching, RoBERTa liu2019roberta, and SimCSE gao-etal-2021-simcse, though not specifically represented in Table LABEL:tab:Classification, remain viable alternatives for developing effective biomedical RAG applications.
@@
AskFDALabel llm-regulatory-framework Sentence Transformer thakur2020augmented FDA drug labeling extraction 2024
CaLM parmanto2024calm BGE-embedding bge_embedding with Chroma DB Supporting caregivers FM 2024
RAG-HPO garcia2024deepphenotyping FastEmbed fastembed Rare genetic disorders automated deep phenotyping 2024
RALL Guo2024 Dense Passage Retriever Karpukhin2020 Lay language generation 2024
RT li2024rt BERT embeddings devlin2019bert Few-shot medical NER 2024
RAG-GPT xu2024evaluation BGE-embedding bge_embedding Breast cancer nursing care QA 2024
GNQA darnell2024biomedical HNSW graphs malkov2016efficient Medical QA 2024
Klang et al. klang2024assessing GIST Large Embedding solatorio2024gistembed with FAISS ICD-10-CM coding 2024
RAGPR zheng2025integrating SBERT embeddings santander2022semantic Personalized physician recommendations 2025
JGCLLM Fukushima2025 GLuCoSE-base-ja vectors Fukuchi2023 Genetic counseling support 2025